# Relatively big projective modules and their applications to direct sum decompositions

### v1

@@ 2. The monoid of relatively big projective modules
Proof.  The isomorphism in $(i)$ is given by $\varphi([X],I)=\iota_{I}([X])$, where $\iota_{I}$ is the embedding from Lemma 2.5. To prove $(ii)$, note that elements in $V(\Lambda)$ can be thought of as elements in $V(\Lambda/I)$ for $I=\{0\}$, so $(ii)$ follows from $(i)$.
The following result gives a characterization of when every element in $B(\Lambda)$ is in the same isomorphism class as a direct sum of finitely generated projective modules.
Proposition 2.8.
Let $\Lambda$ be a ring. Then $B(\Lambda)$ is a semigroup of modules such that any representative is isomorphic to a direct sum of finitely generated projective modules if and only if the following two properties are satisfied:
  1. (1)
    If $I$ is the trace ideal of a projective right $\Lambda$-module, then there is a right $\Lambda$-module $P$, which is a direct sum of finitely generated projective modules such that $\operatorname{Tr}_{\Lambda}(P)=I$.
  2. (2)
    If $I$ is the trace ideal of a countably generated projective right $\Lambda$-module and $P^{\prime}$ is a finitely generated projective right $\Lambda/I$-module, then there is a finitely generated projective right $\Lambda$-module $X$ such that $P^{\prime}\cong X/XI$.
Proof.  Assume that (1) and (2) are satisfied, and let $[P]\in B(\Lambda)$. Then there exists a countably generated projective right $\Lambda$-module $Q$ with trace ideal $J$ such that $P$ is $J$-big and $P/PJ$ is a finitely generated projective right $\Lambda/J$-module.
By (1), $Q$ can be taken to be a direct sum of finitely generated projective modules. By (2) applied to $P/PJ$, there exists a finitely generated projective right $\Lambda$-module $X$ such that $X/XJ\cong P/PJ$. Let $P^{\prime}=X\oplus Q^{(\omega)}$. Clearly, $[P^{\prime}]\in B(\Lambda)$, and by Lemma 1.14, $[P]=[P^{\prime}]$.
Assume now that every element in $B(\Lambda)$ is in the same isomorphic class as a direct sum of finitely generated projective modules. Let $P$ be a projective right $\Lambda$-module with trace $I$. By [kaplansky], $P=\bigoplus_{i\in J}P_{i}$ is a direct sum of countably generated projective modules. By Example 2.3, $[P_{i}^{(\omega)}]\in B(\Lambda)$. Therefore, $P_{i}^{(\omega)}$ is isomorphic to a direct sum of finitely generated projective modules for each $i\in J$. Hence, $P^{(\omega)}$ is isomorphic to a direct sum of finitely generated projective modules and $I=\operatorname{Tr}_{\Lambda}(P^{(\omega)})$. This gives (1).
To prove (2), let $Q$ be a countably generated projective right $\Lambda$-module with trace $I$, and let $P^{\prime}$ be a finitely generated projective right $\Lambda/I$-module. By Lemma 1.9, since $I$ is the trace of a countably generated projective module, there exists a countably generated projective right $\Lambda$-module $P$ such that $I\subseteq\operatorname{Tr}_{\Lambda}(P)$, $P/PI\cong P^{\prime}$ and $\operatorname{Tr}_{\Lambda}(P)/I=\operatorname{Tr}_{\Lambda/I}(P^{\prime})$.
Now, $[P\oplus Q^{(\omega)}]\in B(\Lambda)$, so it is a direct sum of finitely generated projective modules. Write $P\oplus Q^{(\omega)}=\bigoplus_{i\in K}X_{i}$ with $X_{i}$ a finitely generated projective $\Lambda$-module for each $i\in K$. Since $(\bigoplus_{i\in K}X_{i})/(\bigoplus_{i\in K}X_{i})I\cong P^{\prime}$ is finitely generated, there is a finite subset $K_{0}\subseteq K$ such that $(\bigoplus_{i\in K_{0}}X_{i})/(\bigoplus_{i\in K_{0}}X_{i})I\cong P^{\prime}$, so $X=\bigoplus_{i\in K_{0}}X_{i}$ is the finitely generated projective $\Lambda$-module we were looking for.
@@
Let $\Lambda$ be a ring. It is well known that the functor $(-)^{*}=\mathrm{Hom}_{\Lambda}(-,\Lambda)$ induces a duality between finitely generated projective right $\Lambda$-modules and finitely generated projective left $\Lambda$-modules. So that, $\varphi\colon V_{r}(\Lambda)\to V_{\ell}(\Lambda)$ defined by $\varphi[P]=[P^{*}]$, is an isomorphism of monoids. Moreover, $P$ and $P^{*}$ have the same trace (cf. [andersonfuller, Exercise 22.1]).
In the next example, we show that the canonical monoid isomorphism $V_{r}(\Lambda)\cong V_{\ell}(\Lambda)$ of Remark 2.9, in general, does not extend to an isomorphism between the monoids $V_{r}(\Lambda)\sqcup B_{r}(\Lambda)$ and $V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)$. We will show that it does extend, provided that $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$.
@@
In such $\Lambda$, what happens is that one of the maximal ideals contains a nonzero trace of a countably generated projective right $R$-module. Therefore $\mathcal{T}_{\ell}(\Lambda)=\{0\},\Lambda\subsetneq\mathcal{T}_{r}(\Lambda)$. Then $V_{r}(\Lambda)\sqcup B_{r}(\Lambda)$ and $V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)$ are not isomorphic. This cannot happen if $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$.
Assume $\Lambda$ is a ring such that $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$. In view of the operation defined in Remark 2.6, to define a monoid homomorphism from $V_{r}(\Lambda)\sqcup B_{r}(\Lambda)$ to $V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)$, that extends the isomorphism between the monoids of finitely generated projective modules of Remark 2.9, we have to be careful with the compatibility of the operations defined in both monoids. That is, identifying $[P]\in V_{r}(\Lambda/I)$ with $([P],I)\in V_{r}(\Lambda)\sqcup B_{r}(\Lambda)$ and defining the map by $([P],I)\mapsto([P^{*}],I)$, then we must check that $([P_{1}],I)+([P_{2}],J)$ is mapped to $([P_{1}^{*}],I)+([P_{2}^{*}],J)$.
The following result is a technical lemma that is crucial for our discussion.
@@
$\cong\operatorname{Hom}_{\Lambda/I}(P,\operatorname{Hom}_{\Lambda/(I+J)}(\Lambda/(I+J),\Lambda/(I+J)))$
@@
where in the second isomorphism we have used the Hom-tensor adjunction. Now, we want to prove that $\operatorname{Hom}_{\Lambda/I}(P,\Lambda/(I+J))\cong\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)/J\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)$ as left $\Lambda/I$-modules.
@@
$\ker\varphi$ =$=\{f\in\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)\mid\pi\circ f=0\}$
=$=\{f\in\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)\mid f(x)\in\ker\pi\text{ for every }x\in P\}$
=$=\{f\in\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)\mid f(x)\in(I+J)/I\text{ for every }x\in P\}$
@@
${P}$${(\Lambda/I)^{(A)}}$${(I+J)/I}$${0.}$$\scriptstyle{f}$$\scriptstyle{F}$$\scriptstyle{\pi}$
Since $P$ is finitely generated, the image of $F$ is contained in a finite number of copies of $\Lambda/I$, say $(\Lambda/I)^{(A_{0})}$, with $A_{0}\subseteq A$ finite. Let $\{x_{i}\}_{i\in A}$ be a basis of $(\Lambda/I)^{(A)}$. Then
@@
Conversely, let $F=\sum_{i=1}^{n}(y_{i}+I)\cdot F_{i}\in(I+J)/I\cdot\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)$ for some $y_{1},\dotsc,y_{n}\in J$ and $F_{1},\dotsc,F_{n}\in\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)$. Then regarding the product of elements as composition of homomorphisms, $F\in\operatorname{Hom}_{\Lambda/I}(P,(I+J)/I)$. Therefore, by the First Isomorphism Theorem,
@@
Let $\Lambda$ be a ring such that $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$. Then there is an isomorphism of monoids between $V_{r}(\Lambda)\sqcup B_{r}(\Lambda)$ and $V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)$.
Proof.  Let $\varphi\colon V_{r}(\Lambda)\sqcup B_{r}(\Lambda)\to V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)$ be the map defined by $\varphi([P],I)=([P^{*}],I)$. By Remark 2.9, it is well-defined. By Lemma 2.11, it is a monoid homomorphism since
@@
Finally, by Remark 2.9 it is bijective.
Then Theorem 1.18 together with the fact that, over a noetherian ring, $\Lambda$, $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$ is just the set of idempotent ideals of $\Lambda$ (cf. [whitehead]), immediately yields.
@@
[P2, Theorem 2.12] Let $\Lambda$ be a noetherian ring satisfying condition ()(*)( ∗ ). Then $V_{r}^{*}(\Lambda)=V_{r}(\Lambda)\sqcup B_{r}(\Lambda)\cong V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)=V_{\ell}^{*}(\Lambda)$.
In general, it is an open question whether $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$ for $\Lambda$ a right noetherian ring (cf. Corollary 4.9 and the comments after it for further information around this question).
@@ 3. Lifting projectives modules modulo a right $T$-nilpotent ideal
  1. (1)
    Let $X$ be a right projective $\Lambda/N$-module. Then there exists a projective right $\Lambda$-module $P$ such that $P/PN\cong X$. All projective modules with this property are isomorphic. In addition, $X$ is finitely generated or countably generated if and only if so is $P$.
  2. (2)
@@
Proof.  $(1)$ Let $X$ be a projective right $\Lambda/N$-module. As $X$ is a direct sum of countably generated projective right $\Lambda/N$-modules, we may assume that $X$ is countably generated. In this case, $X$ is the direct limit of a direct system
[…]
@@
[…]
such that when we apply the functor $-\otimes_{\Lambda}\Lambda/N$ to (3) we obtain a direct system isomorphic to (1).
@@
As (2) splits, there exists $\Psi\colon\bigoplus_{n\geq 1}{F}_{n}\to\bigoplus_{n\geq 1}{F}_{n}$ such that $\Psi\circ\Phi=\mathrm{Id}+B$ where $B$ can be thought of as a column finite matrix with entries in $N$ (that is, $B\colon\bigoplus_{n\geq 1}{F}_{n}\to\bigoplus_{n\geq 1}{F}_{n}$ has the image contained in $\bigoplus_{n\geq 1}{F}_{n}N$). Since $N$ is right $T$-nilpotent, $B$ is in the Jacobson radical of the ring of countable column-finite matrices with coefficients in $\Lambda$ [SW], and then $\mathrm{Id}+B$ is invertible. So that $(\mathrm{Id}+B)^{-1}\circ\Psi$ is a left inverse of $\Phi$ and we can deduce that $P$ is projective.
This shows that any lifting of the direct system (1) gives a projective $\Lambda$-module $P$ that, in addition, satisfies $P/PN\cong X$. As $PN$ is a small submodule of $P$ by [andersonfuller, Lemma 28.3], $P/PN\cong X$ is finitely generated (countably generated) if and only if so is $P$.
$(2)$. By $(1)$, the module $X$ can be lifted to a countably generated projective right $\Lambda$-module $P$ such that $P/PN\cong X$. Let $J$ be the trace of $P$, then $J+N=I+N$. By [traces, Corollary 2.9], $J\subseteq I$ because $J$ is a trace ideal of a projective module and $N$, being right $T$-nilpotent, is contained in the Jacobson radical of $\Lambda$. Since $J+N=I+N$ and $I^{2}=I$, $IJ+IN=I$. By [andersonfuller, Lemma 28.3], $IN\ll I$ which implies $IJ=I$ and, in particular, $I\subseteq J$. Therefore, $I=J$ is the trace of $P$.
@@
Proof.  Assume that $P$ is $I$-big. Let $X$ be a countably generated projective right $\Lambda/N$-module such that its trace is contained in $(I+N)/N$. By Lemma 3.1, there exists a countably generated projective right $\Lambda$-module $Q$ such that $Q/QN\cong X$; therefore its trace $J$ is contained in $I+N$. Since $J$ is the trace of a countably generated projective, and $N$ is right $T$-nilpotent, we deduce from [traces, Corollary 2.9] that $J\subseteq I$. By hypothesis, there exists an onto module homomorphism $f\colon P\to Q$. Therefore, there is an onto module homomorphism
${g\colon P}$${Q}$${Q/QN\cong X}$$\scriptstyle{f}$$\scriptstyle{\pi}$
@@
To prove the converse, let $Q$ be a countably generated projective right $\Lambda$-module such that its trace is contained in $I$. Then the trace of $Q/QN$ is contained in $I+N/N$ and since $P/PN$ is $(I+N)/N$-big, there exists an onto module homomorphism $g\colon P/PN\to Q/QN$. Since $P$ is projective, there exists $f\colon P\to Q$ such that the diagram
${P}$${Q}$${Q/QN}$$\scriptstyle{f}$$\scriptstyle{g\circ\pi_{1}}$$\scriptstyle{\pi_{2}}$
@@
Let $\Lambda$ be a ring, and let $N$ be a right $T$-nilpotent ideal of $\Lambda$. Then the functor $-\otimes_{\Lambda}\Lambda/N$ induces a monoid isomorphism $\varphi\colon V^{*}(\Lambda)\to V^{*}(\Lambda/N)$ which restricts to monoid isomorphism $V(\Lambda)\to V(\Lambda/N)$ and $V(\Lambda)\sqcup B(\Lambda)\to V(\Lambda/N)\sqcup B(\Lambda/N)$.
Proof.  Consider the monoid morphism $\varphi\colon V^{*}(\Lambda)\to V^{*}(\Lambda/N)$ defined by $\varphi([P])=[P/PN]$ where $P$ denotes a countably generated, projective, right $\Lambda$-module. It follows from Lemma 3.1 that $\varphi$ is an isomorphism that restricts to an isomorphism $V(\Lambda)\to V(\Lambda/N)$.
Let $P$ be a countably generated projective $\Lambda$-module such that $[P]\in B(\Lambda)$. That is there exists $I$, the trace of a nonzero countably generated projective module $Q$, such that $P$ is $I$-big and $P/PI$ is finitely generated.
By Lemma 1.6, $I+N/N$ is the trace of $Q/QN$. Notice also that $I+N/N\neq\{0\}$. By Lemma 3.2, $P/PN$ is $I+N/N$-big. Clearly
@@
is finitely generated. Hence $\varphi([P])\in B(\Lambda/N)$.
Let $X$, $Y$ be countably generated projective right $\Lambda/N$-modules. Let $K$ be the trace of $Y$ in $\Lambda/N$. Assume that $X$ is $K$-big and that $X/XK$ is finitely generated.
By Lemma 3.1 there exist countably generated projective $\Lambda$-modules $P$ and $Q$ such that $P/PN\cong X$ and $Q/QN\cong Y$. Let $I$ be the trace of $Q$. By Lemma 1.6, $K=I+N/N$.
By Lemma 3.2, $P$ is $I$-big. Moreover, $P/PI$ is finitely generated because so is $P/P(I+N)$. This shows that $\varphi^{-1}([X])\in B(\Lambda)$. This concludes the proof that $\varphi$ induces an isomorphism $V(\Lambda)\sqcup B(\Lambda)\to V(\Lambda/N)\sqcup B(\Lambda/N)$.

### v2

@@ 2. The monoid of relatively big projective modules
Proof. The isomorphism in $(i)$ is given by $\iota([X],I)=\iota_{I}([X])$, where $\iota_{I}$ is the embedding from Lemma 2.3. By Lemma 1.14, $\iota$ is bijective.
To prove $(ii)$, note that elements in $V(\Lambda)$ can be thought of as elements in $V(\Lambda/I)$ for $I=\{0\}$, so $(ii)$ follows from $(i)$.
@@
Let $\Lambda$ be a ring. It is well known that the functor $(-)^{*}=\mathrm{Hom}_{\Lambda}(-,\Lambda)$ induces a duality between finitely generated projective right $\Lambda$-modules and finitely generated projective left $\Lambda$-modules. So that, $\varphi\colon V_{r}(\Lambda)\to V_{\ell}(\Lambda)$ defined by $\varphi([P])=[P^{*}]$, is an isomorphism of monoids. Moreover, $P$ and $P^{*}$ have the same trace (cf. [andersonfuller, Exercise 22.1]).
In the next example, we show that the canonical monoid isomorphism $V_{r}(\Lambda)\cong V_{\ell}(\Lambda)$ of Remark 2.6 does not, in general, extend to an isomorphism between the monoids $V_{r}(\Lambda)\sqcup B_{r}(\Lambda)$ and $V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)$. We will show that it does extend, provided that $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$.
@@
In such $\Lambda$, what happens is that one of the maximal ideals contains a non-zero trace of a countably generated projective right $R$-module. Therefore $\mathcal{T}_{\ell}(\Lambda)=\{0,\Lambda\}\subsetneq\mathcal{T}_{r}(\Lambda)$. Then $V_{r}(\Lambda)\sqcup B_{r}(\Lambda)$ and $V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)$ are not isomorphic. This cannot happen if $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$.
Assume $\Lambda$ is a ring such that $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$. We define a map $V_{r}(\Lambda)\sqcup B_{r}(\Lambda)\mapsto V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)$ that extends the isomorphism between the monoids of finitely generated projective modules of Remark 2.6 by $([P],I)\mapsto([P^{*}],I)$ for any $([P],I)\in V_{r}(\Lambda)\sqcup B_{r}(\Lambda)$. We must check that $([P_{1}],I)+([P_{2}],J)$ is mapped to $([P_{1}^{*}],I)+([P_{2}^{*}],J)$.
The following technical lemma is crucial for our discussion.
@@
where in the second isomorphism, we have used the Hom-tensor adjunction. Now, we want to prove that $\operatorname{Hom}_{\Lambda/I}(P,\Lambda/(I+J))\cong\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)/J\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)$ as left $\Lambda/I$-modules.
@@
$\ker\varphi$ $=\{f\in\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)\mid\pi\circ f=0\}$
$=\{f\in\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)\mid f(x)\in(I+J)/I\text{ for every }x\in P\}$
@@
[…]
Since $P$ is finitely generated, the image of $F$ is contained in a finite number of copies of $\Lambda/I$, namely $(\Lambda/I)^{(A_{0})}$, with $A_{0}\subseteq A$ finite. Let $\{x_{i}\}_{i\in A}$ be the canonical basis of $(\Lambda/I)^{(A)}$. Then
@@
Conversely, let $F=\sum_{i=1}^{n}(y_{i}+I)\cdot F_{i}\in(I+J)/I\cdot\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)$ for some $y_{1},\dotsc,y_{n}\in J$ and $F_{1},\dotsc,F_{n}\in\operatorname{Hom}_{\Lambda/I}(P,\Lambda/I)$. Then, regarding the product of elements as a composition of homomorphisms, $F\in\operatorname{Hom}_{\Lambda/I}(P,(I+J)/I)$. Therefore, by the First Isomorphism Theorem,
@@
Let $\Lambda$ be a ring such that $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda):=\mathcal{T}(\Lambda)$. Then there is an isomorphism of monoids between $V_{r}(\Lambda)\sqcup B_{r}(\Lambda)$ and $V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)$.
Proof. Let $\varphi\colon\bigsqcup_{I\in\mathcal{T}(\Lambda)}V_{r}(\Lambda/I)\to\bigsqcup_{I\in\mathcal{T}(\Lambda)}V_{\ell}(\Lambda/I)$ be the map defined by $\varphi([P],I)=([P^{*}],I)$. By Remark 2.6, it is well-defined and bijective. By Lemma 2.8, it is a monoid homomorphism since
@@
Finally, as $\varphi$ is a monoid isomorphism, by Corollary 2.5 there is also a monoid isomorphism $V_{r}(\Lambda)\sqcup B_{r}(\Lambda)\to V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)$ .
Then, Theorem 1.23, together with the fact that over a noetherian ring $\Lambda$, $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$ is just the set of idempotent ideals of $\Lambda$ (cf. [whitehead]), immediately yields.
@@
[P2, Theorem 2.12] Let $\Lambda$ be a noetherian ring satisfying condition ()(*). Then $V_{r}^{*}(\Lambda)=V_{r}(\Lambda)\sqcup B_{r}(\Lambda)\cong V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)=V_{\ell}^{*}(\Lambda)$.
In general, it is an open question whether $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$ for $\Lambda$ a right noetherian ring (cf. Corollary 4.9 and the comments after it for further information regarding this question).
@@ 3. Lifting projectives modules modulo a right $T$-nilpotent ideal
  1. (i)
    Let $X$ be a projective right $\Lambda/N$-module. Then there exists a projective right $\Lambda$-module $P$ such that $P/PN\cong X$. All projective modules with this property are isomorphic. In addition, $X$ is finitely generated (countably generated) if and only if $P$ is finitely generated (countably generated).
  2. (ii)
    Let $P$ and $Q$ be projective right $\Lambda$-modules. Let $f\colon P\to Q$ be a morphism of $\Lambda$-modules. Then $f$ is an isomorphism if and only if the induced map $P/PN\to Q/QN$ is an isomorphism of $\Lambda/N$-modules.
  3. (iii)
@@
Proof. $(i)$ Let $X$ be a projective right $\Lambda/N$-module. As $X$ is a direct sum of countably generated projective right $\Lambda/N$-modules, we may assume that $X$ is countably generated. In this case, $X$ is the direct limit of a direct system
$\overline{F}_{1}$$\overline{F}_{2}$$\cdots$$\overline{F}_{n}$$\cdots$} (1)
@@
$F_{1}$$F_{2}$$\cdots$$F_{n}$$\cdots$} (3)
such that when we apply the functor $\otimes_{\Lambda}\Lambda/N$ to (3), we obtain a direct system isomorphic to (1).
@@
As (2) splits, there exists $\Psi\colon\bigoplus_{n\geq 1}{F}_{n}\to\bigoplus_{n\geq 1}{F}_{n}$ such that $\Psi\circ\Phi=\mathrm{Id}+B$ where $B$ can be thought of as a column finite matrix with entries in $N$ (that is, $B\colon\bigoplus_{n\geq 1}{F}_{n}\to\bigoplus_{n\geq 1}{F}_{n}$ has its image contained in $\bigoplus_{n\geq 1}{F}_{n}N$). Since $N$ is right $T$-nilpotent, $B$ is in the Jacobson radical of the ring of the $\mathbb{N}\times\mathbb{N}$ column-finite matrices with coefficients in $\Lambda$ [SW], and therefore, $\mathrm{Id}+B$ is invertible. So that $(\mathrm{Id}+B)^{-1}\circ\Psi$ is a left inverse of $\Phi$, and we can deduce that $P$ is projective.
This shows that any lifting of the direct system (1) gives a projective $\Lambda$-module $P$ that, in addition, satisfies $P/PN\cong X$. As $PN$ is a small submodule of $P$ by [andersonfuller, Lemma 28.3], $P/PN\cong X$ is finitely generated (countably generated) if and only if $P$ is also finitely generated (countably generated).
The uniqueness, up to isomorphism, of $P$ follows from $(ii)$.
$(ii)$ If $f\colon P\to Q$ is an isomorphism of $\Lambda$-modules, then it has an inverse homomorphism $g\colon Q\to P$. Let $\overline{f}\colon P/PN\to Q/QN$ and $\overline{g}\colon Q/QN\to P/PN$ be the induced maps. Then $\overline{g}$ is the inverse of $\overline{f}$, so both are also invertible morphisms of $\Lambda/N$-modules.
Conversely, if $\alpha\colon P/PN\to Q/QN$ is an isomorphism of $\Lambda/N$-modules. As $PN$ is a small submodule of $P$ by [andersonfuller, Lemma 28.3], $P$ and $Q$ are projective covers of $P/PN$. Therefore, any lifting of $\alpha$ to a morphism $P\to Q$ is an isomorphism.
$(iii)$. By $(i)$, the module $X$ can be lifted to a countably generated projective right $\Lambda$-module $P$ such that $P/PN\cong X$. Let $J$ be the trace of $P$. Then, by Lemma 1.6, $J+N=I+N$. By [traces, Corollary 2.9], $J\subseteq I$ because $J$ is a trace ideal of a projective module, and $N$, being right $T$-nilpotent, is contained in the Jacobson radical of $\Lambda$. Since $J+N=I+N$ and $I^{2}=I$, $IJ+IN=I$. By [andersonfuller, Lemma 28.3], $IN\ll I$, which implies $IJ=I$ and, in particular, $I\subseteq J$. Therefore, $I=J$ is the trace of $P$.
@@
Proof. Assume that $P$ is $I$-big. Let $X$ be a countably generated projective right $\Lambda/N$-module such that its trace is contained in $(I+N)/N$. By Lemma 3.1, there exists a countably generated projective right $\Lambda$-module $Q$ such that $Q/QN\cong X$. Therefore, its trace $J$ is contained in $I+N$. Since $J$ is the trace of a countably generated projective, and $N$ is right $T$-nilpotent, we deduce from [traces, Corollary 2.9] that $J\subseteq I$. By hypothesis, there exists an onto module homomorphism $f\colon P\to Q$. Therefore, there is an onto module homomorphism
[…]
@@
To prove the converse, let $Q$ be a countably generated projective right $\Lambda$-module such that its trace is contained in $I$. By Lemma 1.6, the trace of $Q/QN$ is contained in $I+N/N$, and since $P/PN$ is $(I+N)/N$-big, there exists an onto module homomorphism $g\colon P/PN\to Q/QN$. Since $P$ is projective, there exists $f\colon P\to Q$ such that the diagram
[…]
@@
Let $\Lambda$ be a ring, and let $N$ be a right $T$-nilpotent ideal of $\Lambda$. Then the functor $\otimes_{\Lambda}\Lambda/N$ induces a monoid isomorphism $\varphi\colon V^{*}(\Lambda)\to V^{*}(\Lambda/N)$ which restricts to monoid isomorphisms $V(\Lambda)\to V(\Lambda/N)$ and $V(\Lambda)\sqcup B(\Lambda)\to V(\Lambda/N)\sqcup B(\Lambda/N)$.
Proof. Consider the monoid morphism $\varphi\colon V^{*}(\Lambda)\to V^{*}(\Lambda/N)$ defined by $\varphi([P])=[P/PN]$, where $P$ denotes a countably generated, projective, right $\Lambda$-module. It follows from Lemma 3.1 that $\varphi$ is an isomorphism. More precisely, $\varphi$ is injective by the claim $(2)$ of Lemma 3.1, and it is onto by claim $(1)$. Claim $(1)$ of Lemma 3.1 also implies that $\varphi$ restricts to an isomorphism $V(\Lambda)\to V(\Lambda/N)$.
It remains to prove that $\varphi$ restricts to an isomorphism $B(\Lambda)\to B(\Lambda/N)$.
Let $P$ be a countably generated projective $\Lambda$-module such that $P]\in B(\Lambda)$. That is, there exists $I$, the trace of a non-zero countably generated projective module $Q$, such that $P$ is $I$-big and $P/PI$ is finitely generated.
By Lemma 1.6, $(I+N)/N$ is the trace of $Q/QN$. Notice also that $(I+N)/N\neq\{0\}$. By Lemma 3.2, $P/PN$ is $(I+N)/N$-big. Clearly,
@@
is finitely generated. Hence, $\varphi([P])\in B(\Lambda/N)$.
Let $X$ be a countably generated projective right $\Lambda/N$-module such that $X]\in B(\Lambda/N)$. This means that there exists a non-zero countably generated projective right $\Lambda/N$-module with trace ideal $K$, such that $X$ is $K$-big and $X/XK$ is finitely generated.
By Lemma 3.1, there exist countably generated projective $\Lambda$-modules $P$ and $Q$ such that $P/PN\cong X$ and $Q/QN\cong Y$. Let $I$ be the trace of $Q$. By Lemma 1.6, $K=(I+N)/N$.
By Lemma 3.2, $P$ is $I$-big. Moreover, $P/PI$ is finitely generated because $P/P(I+N)$ is also finitely generated. This shows that $P]=\varphi^{-1}([X])\in B(\Lambda)$.
This concludes the proof that $\varphi$ induces an isomorphism $B(\Lambda)\to B(\Lambda/N)$.